Problem 3 a) Find the Present Value at Year 0 (5 points) b) Find the Future Value at Year 10 (5 points) $700 $600A $500 ▲ $400 ▲ 5% $300 $200 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14
Answer #1

ANSWER:

A) Present value:

i = 5%

pv = cash flow in year 5(p/f,i,n) + cash flow in year 6(p/f,i,n) + cash flow in year 7(p/f,i,n) + cash flow in year 8(p/f,i,n) + cash flow in year 9(p/f,i,n) + cash flow in year 10(p/f,i,n)

pv = 200(p/f,5%,5) + 300(p/f,5%,6) + 400(p/f,5%,7) + 500(p/f,5%,8) + 600(p/f,5%,9) + 700(p/f,5%,10)

pv = 200 * 0.7835 + 300 * 0.7462 + 400 * 0.7107 + 500 * 0.6768 + 600 * 0.6446 + 700 * 0.6139

pv = 156.7 + 223.86 + 284.28 + 338.4 + 386.76 + 429.73

pv = 1,819.73

so the present value is $1,819.73

B) Future value:

i = 5% and n = 10 years

fv = pv(f/p,i,n)

fv = 1,819.73(f/p,5%,10)

fv = 1,819.73 * 1.629

fv = 2,964.34

so the future value is 2,964.34